OpenStudy (anonymous):

Factoring questions? Determine which answer choice below is equivalent to the given expression by completing the square.

OpenStudy (anonymous):

1/3x^2+2x-3

OpenStudy (anonymous):

Here's how I would simplify:\[\frac{ 1 }{ 3}x^2+2x-3=\frac{ 1 }{ 3 }[x^2+\frac{ 2 }{ 3}-1]=\frac{ 1 }{ 3 }[(x+\frac{ 1 }{ 3 })^2-\frac{ 1 }{ 9 }-1]=\frac{ 1 }{ 3 }[(x+\frac{ 1 }{ 3 })^2-\frac{ 10 }{ 9 }]\]

OpenStudy (anonymous):

omg thank you so much! this is so confusing, my teacher didnt do a good job of explaining